ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 46 Lee, Sa Ge, Seoul National University, Korea The generalized normal state space and unital normal completely positive map By introducing the notion of a generalized normal state space, we give a necessary and sufficient condition for that there exists a unital normal completely positive map from a von Neumann algebra into another, in terms of their generalized normal state spaces. Reference: Sa Ge Lee, The generalized normal state space and unital normal completely positive map, J. Korean Math. Soc.35(1998), No.1, 237-257.